I am having troublw with this question:
Find the radius and center of x^2+y^2+12x=0


Thanks,

Answer by fractalier(1829)

Okay, we will need to complete the square on the x's...so from
x^2 + y^2 + 12x = 0
Let us rearrange and complete the square
x^2 + 12x + ___ + y^2 = 0 + ___
What goes in the blank is half of 12, squared, or 36
x^2 + 12x + 36 + y^2 = 0 + 36
Now we have
(x + 6)^2 + y^2 = 36
which is a circle centered at (-6, 0) and radius of 6